OpenStudy (anonymous):

solve the system of equations using any method? 2x+5y=14 6x+7y=10

OpenStudy (anonymous):

what you have learned from previous problem? ok lets solve this together

OpenStudy (anonymous):

subtract 5y from both sides of equation 1 tell me that you get.

OpenStudy (anonymous):

9

OpenStudy (anonymous):

try again !!!

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

ok lets do this subtract 5y from both sides 2x+5y-5y=14-5y now +5y and -5y will get cancelled on the left side. it becomes 2x=14-5y is it ok so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok divide both sides by 2 (2x)/(2)=(14-5y)/2 \[\Large \implies x=\frac{14-5y}{2}\] now ?

OpenStudy (anonymous):

9y divided by 2

OpenStudy (anonymous):

how 9y divide /2 can you explain? it is 14 not 14y that you are subtracting.

OpenStudy (anonymous):

oh okay!! now whats next

OpenStudy (anonymous):

idk how to do this!! sorry!!

OpenStudy (anonymous):

ok there was typo in my response it is 6 not 3 in the second equation the coefficient of x. correct is below \[\Large \implies 6((\frac{14-5y}{2})+7y=10\] six can be divided by 2 to get 3 \[\Large \implies 3(14-5y)+7y=10\] now can you simplify?

OpenStudy (anonymous):

42-15y+7y=10

OpenStudy (anonymous):

?? is this coreect?

OpenStudy (anonymous):

good!!

OpenStudy (anonymous):

now -15y+7y =???

OpenStudy (anonymous):

22y

OpenStudy (anonymous):

try again !!

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

correct!!!

OpenStudy (anonymous):

now the equation becomes 42-8y=10 solve for y. (first move 42 to the right sides, then divide by -8)

OpenStudy (anonymous):

y=-4

OpenStudy (anonymous):

?? is this correct??

OpenStudy (anonymous):

it should be +4 not -4

OpenStudy (anonymous):

because -8y=-32 y=4

OpenStudy (anonymous):

now put this in any of the given equations to get x

hero (hero):

Multiply the first equation by 3, then set 6x = 6x

OpenStudy (anonymous):

@broderick365 did you find x?

OpenStudy (anonymous):

2(4)+5y=14 8x+5y=14 -5y --5 8x =9 this is where i stopped

OpenStudy (anonymous):

?? @sami-21

OpenStudy (anonymous):

use y=4 ! 2x+5(4)=14 now find x.

OpenStudy (anonymous):

2x+5(4)=14 2x+20=14 -20 -20 2x =34 is th right so far??

OpenStudy (anonymous):

@sami-21 ??

OpenStudy (anonymous):

no it is not. try again. it is correct till 2x+20=14 move from here try subtracting 20 from both sides then divide by 2

OpenStudy (anonymous):

i did i got -5

OpenStudy (anonymous):

2x+20-20=14-20 +20 and -20 get cancelled on the left side you are left with 2x=14-20 now try.

OpenStudy (anonymous):

2x=5

OpenStudy (anonymous):

14-20 ????

OpenStudy (anonymous):

ohhhhhhh 6 my bad -6!!

OpenStudy (anonymous):

ok now you have 2x=-6 so what is x ?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

yes